# Brain network science modelling of sparse neural networks enables Transformers and LLMs to perform as fully connected

### v3

@@ Appendix H Ablation and Sensitivity Tests
For sparse topological initialization, we compare BRF, BSW, BSF, and CSTI [zhang2023ultra] across three image classification datasets, as shown in Figure 8b. The results indicate that when the inputs can directly access task-relevant information, CSTI consistently achieves the best performance. In general, BRF and BSW perform similarly under these conditions, but outperform the BSF initialization.
@@
We first conduct a simple evaluation of the link removal methods introduced in this article when changing the $\alpha$ and $\delta$ inside Figure 1b2) on Figure 6b) and Figure 7b). The removal methods are selected from Weight Magnitude (WM), Weight Magnitude soft (WMs), Relative Importance (RI), and Relative Importance soft (RIs). For WM we fix the hyperparameters $\alpha=1$ and $\delta=1$; for RI we fix the hyperparameters $\alpha=0$ and $\delta=0.5$; for WMs we fix $\alpha=1$ and we let $\delta$ increase linearly from $0.5$ to $0.9$; for RIs we fix $\alpha=0$ and let $\delta$ increase linearly from $0.5$ to $0.9$. From the results, it can be observed that WMs performs the best in most cases. We compare these methods with those in [nowak2023fantasticweightsthemprune] in Table 12 on two machine translation tasks. The results indicate that using WMs as a link removal method generally outperforms the alternatives.
@@
To further demonstrate the benefit of the density decay strategy, we conduct an ablation study comparing the sigmoid-based density decay strategy with a cubic-based density decay strategy. We evaluate the performance on the LLaMA-60M model using the C4 dataset and report the perplexity. We show the number of density decay steps and a comparison between the cubic density decay that GraNet [liu2021granet] and GMP [zhu2017gmp] introduced and our proposed sigmoid density decay strategy. The results in the Table 13 show that the sigmoid-based decay consistently achieves lower PPL than the cubic-based decay in all the density decay steps, and both strategies outperform the baseline fixed sparsity (CHTs). These results confirm the motivation for introducing the sigmoid density decay strategy. The table below shows the PPL of CHTss with the two density decay strategies over different density decay steps.
@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
In this section, we show the formula for the proposed Sigmoid Density Decay and the Cubic decay implemented in GraNet [liu2021granet] and GMP [han2015gmp].
@@
Inspired by GMP [han2015gmp, zhu2017gmp], we incorporate historical weights into our CHTs and CHTss implementation. During training, we maintain a historical weight matrix that records previously learned weights throughout the training process. When CHTs and CHTss predict new links, we initialize them using their corresponding historical weights - specifically, the values they held before being pruned. In this way, CHTs and CHTss enable weight recovery with preserved memory, allowing the model to retain valuable prior information.
@@
Table 20: Zero-shot evaluation of LLaMA-1B across GLUE and SuperGLUE. ACC scores are shown. Values are mean ± sd over 5 seeds (lm-eval [eval-harness]). Red values indicate the top scores over models and sparsities for a benchmark.
@@
More recently, BiDST has elegantly reframed DST as a bi-level optimization problem to co-optimize the weights and the sparsity mask simultaneously [pmlr-v235-ji24a]. Other methods improve the training dynamics itself, such as AC/DC, which alternates between sparse and dense phases [peste2021acdc], or Powerpropagation, which introduces a sparsity-inducing weight reparameterization [schwarz2021powerpropagation]. To enhance generalization in the often-chaotic loss landscape of sparse models, S2-SAM provides an efficient, plug-and-play sharpness-aware optimizer [ji2024s2sam]. Another thrust of research adapts DST for specific architectural or hardware advantages. For example, SLaK utilizes dynamic sparsity to enable massive, high-performance kernels in CNNs [liu2022slak]. Chase provides a practical method to translate unstructured dynamic sparsity into hardware-friendly, channel-level sparsity [yin2023chase]. Addressing network health, NeurRev proposes a mechanism to revitalize “dormant neurons” that can emerge during training [li2024neurrev].

### v4

@@ Appendix H Ablation and Sensitivity Tests
For sparse topological initialization, we compare BRF, BSW, BSF, and CSTI [61] across three image classification datasets, as shown in Figure 8b. The results indicate that when the inputs can directly access task-relevant information, CSTI consistently achieves the best performance. In general, BRF and BSW perform similarly under these conditions, but outperform the BSF initialization.
@@
We first conduct a simple evaluation of the link removal methods introduced in this article when changing the $\alpha$ and $\delta$ inside Figure 1b2) on Figure 6b) and Figure 7b). The removal methods are selected from Weight Magnitude (WM), Weight Magnitude soft (WMs), Relative Importance (RI), and Relative Importance soft (RIs). For WM we fix the hyperparameters $\alpha=1$ and $\delta=1$; for RI we fix the hyperparameters $\alpha=0$ and $\delta=0.5$; for WMs we fix $\alpha=1$ and we let $\delta$ increase linearly from $0.5$ to $0.9$; for RIs we fix $\alpha=0$ and let $\delta$ increase linearly from $0.5$ to $0.9$. From the results, it can be observed that WMs performs the best in most cases. We compare these methods with those in [41] in Table 12 on two machine translation tasks. The results indicate that using WMs as a link removal method generally outperforms the alternatives.
@@
To further demonstrate the benefit of the density decay strategy, we conduct an ablation study comparing the sigmoid-based density decay strategy with a cubic-based density decay strategy. We evaluate the performance on the LLaMA-60M model using the C4 dataset and report the perplexity. We show the number of density decay steps and a comparison between the cubic density decay that GraNet [35] and GMP [63] introduced and our proposed sigmoid density decay strategy. The results in the Table 13 show that the sigmoid-based decay consistently achieves lower PPL than the cubic-based decay in all the density decay steps, and both strategies outperform the baseline fixed sparsity (CHTs). These results confirm the motivation for introducing the sigmoid density decay strategy. The table below shows the PPL of CHTss with the two density decay strategies over different density decay steps.
@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
In this section, we show the formula for the proposed Sigmoid Density Decay and the Cubic decay implemented in GraNet [35] and GMP [21].
@@
Inspired by GMP [21, 63], we incorporate historical weights into our CHTs and CHTss implementation. During training, we maintain a historical weight matrix that records previously learned weights throughout the training process. When CHTs and CHTss predict new links, we initialize them using their corresponding historical weights - specifically, the values they held before being pruned. In this way, CHTs and CHTss enable weight recovery with preserved memory, allowing the model to retain valuable prior information.
@@
Table 20: Zero-shot evaluation of LLaMA-1B across GLUE and SuperGLUE. ACC scores are shown. Values are mean ± sd over 5 seeds (lm-eval [19]). Red values indicate the top scores over models and sparsities for a benchmark.
@@
More recently, BiDST has elegantly reframed DST as a bi-level optimization problem to co-optimize the weights and the sparsity mask simultaneously [26]. Other methods improve the training dynamics itself, such as AC/DC, which alternates between sparse and dense phases [42], or Powerpropagation, which introduces a sparsity-inducing weight reparameterization [44]. To enhance generalization in the often-chaotic loss landscape of sparse models, S2-SAM provides an efficient, plug-and-play sharpness-aware optimizer [25]. Another thrust of research adapts DST for specific architectural or hardware advantages. For example, SLaK utilizes dynamic sparsity to enable massive, high-performance kernels in CNNs [36]. Chase provides a practical method to translate unstructured dynamic sparsity into hardware-friendly, channel-level sparsity [56]. Addressing network health, NeurRev proposes a mechanism to revitalize “dormant neurons” that can emerge during training [32].